OpenStudy (anonymous):

A line segment has endpoints A(4, 8) and B(2, 10). The point M is the midpoint of AB. What is an equation of a line perpendicular to AB and passing through M?

OpenStudy (anonymous):

ok you need to find the slope of the line

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

use this formula \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } = m\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got y=-x+4

OpenStudy (anonymous):

ok m is -1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

instead of 4 would it be

OpenStudy (anonymous):

y=-x+c

OpenStudy (anonymous):

plug in 1 of the point to the equation like this 8 = -1(4) +c Solve for c

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

i got -4

OpenStudy (anonymous):

umm c should be 12 8 = -4 +12

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now we find the midpoint of this line with this formula \[(\frac{ x _{1}+x _{2} }{ 2 } , \frac{ y _{1}+y _{2} }{ 2 })\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

your midpoint should be (3,9)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a. y = x + 4 b. y = −x + 4 c. y = −x + 6 d. y = x + 6 this is what I have to chose from

OpenStudy (anonymous):

gradient of perpendicular line is \[m _{1}m _{1} = -1\] Since m1 is -1, we can assume m2 is 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so now we have y=x+c

OpenStudy (anonymous):

we plug in the points of the midpoint into the equation we get 9 = 3 + c

OpenStudy (anonymous):

it should be easy from here

OpenStudy (anonymous):

12

OpenStudy (anonymous):

umm no...look at it this way...3+ what number gives you 9?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yep so there you have it y=x+6

OpenStudy (anonymous):

N is the midpoint of segment MP. O is the midpoint of segment NP. Segment MP with midpoint N and O M is located at (7, 8), and P is located at (14, -3). What are the coordinates of O? Select one: a. (-8.75, 4.25) b. (24.50, -0.50) c. (5.25, 1.25) d. (12.25, -0.25)

OpenStudy (anonymous):

ok looks confusing but this is what they want

OpenStudy (anonymous):

Find the midpoint of MP with the formula i gave you earlier

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

would it b C?

OpenStudy (anonymous):

I meant to say would it be C

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

you want the solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK midpoint of MP is \[(\frac{ 7+14 }{ 2 },\frac{ 8+(-3) }{ 2 }) = (10.5, 2.5)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

N = (10.5,2.5), O is the midpoint of NP so \[(\frac{ 10.5+14 }{ 2 }, \frac{ 2.5+(-3) }{ 2 }) = (12.25,-0.25)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

A triangle has vertices at (4, 5), (4, -4), and (-4, -3). What is the approximate perimeter of the triangle? Select one: a. 27 b. 29 c. 28 d. 30 I got c for this on I just wanted to make sure if it was right or wrong

OpenStudy (anonymous):

ok you need to find distance between the points with this formula \[\sqrt{(x _{2}-x _{1})+(y _{2}-y _{1})}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so would I input which ever number to find the answer

OpenStudy (anonymous):

no you need to get 2 points and put it into the formula to get the distance of 1 side of the triangle

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

would be easier if you label it like A = (4,5) B = (4,-4) C =(-4,-3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Find the distance of AB and AC and BC then add them up

OpenStudy (anonymous):

25?

OpenStudy (anonymous):

lemme try

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I got 27

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ill give you the distance you try and work it out AB = 9 AC = 11.31 BC = 8.06

OpenStudy (anonymous):

I got 28 as well :D

OpenStudy (anonymous):

i mean use the formula and try to see if you get the same distance

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The volume of a sphere is 63089.813626667 cubic inches. What is the approximate diameter of this sphere? (Volume of a sphere = \frac{4}{3}\pi {r^3}) Select one: a. 5023.1 in b. 203.4 in c. 49.4 in d. 24.7 in I got b for this and I worked it out but I'm still not sure

OpenStudy (anonymous):

\[63089.813626667 = \frac{ 4 }{ 3 } \pi r^3\] \[63089.813626667(\frac{ 3 }{ 4 }) = \pi r^3\] \[47317.36022 = \pi r^3\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ 47317.36022 }{ \pi } = 15061.58355 = r^3\] \[\sqrt[3]{15061.58355} = 24.69582535 = r\] \[dia. = 2r =24.69582535\times2 =49.39165069\]

OpenStudy (anonymous):

so it's C

OpenStudy (anonymous):

dayum sneaky teacher putting radius as an option